The length of a rectangle is three times its width.
If the perimeter of the rectangle is 72 in, find its area. How would I go about solving this?

Answer by checkley77(12844)

L=2W
P=2L+2W
72=2*3W+2W
72=6W+2W
72=8W
8W=72
W=72/8
W=9 IN. IS THE WIDTH.
L=3*9=27 IN. IS THE LENGTH.pROOF:
72=2*9+2*27
72=18+54
72=72